A rhinoceros is at the origin of coordinates at t1 = 0. For the time interval from 0 to 12 sec, the rhino’s average velocity has x-component -3.8 m/s and y-component 4.9 m/s. At t2 = 12 sec, (a) what are the x- and y-component of the rhino’s position? (b) How far is the rhino from the origin?


1
Expert's answer
2021-07-12T12:12:04-0400

a)


"x=-3.8(12)=-45.6\ m\\
y=4.9(12)=58.8\ m\\"

b)


"d=\sqrt{45.6^2+58.8^2}=74
.4\ m"